DeAndre Hunter Enterprises has found that its common equity capital shares have a beta equal to 1.5 while the risk-free return is 8 percent and the expected return on the market is 14 percent.   The firm is financed with $120,000,000 of common shares (market value) and $80,000,000 of debt, and 35 percent is the marginal tax rate.

The firm has semiannual bonds outstanding with nine years to maturity and are currently priced at $754.08. The bonds have a coupon rate of 7.25 percent. The face value on a bond is $1,000. The firm will invest $300,000 in a five-year project today that will produce annual cash flows of $80,000.

What is the WACC for the firm?
Answer #1

cost of equity =risk free rate+beta*(market return-risk free rate)

=8+1.5*(14-8) =17%

---

cost of debt =YTM of bond => can be calculated in excel as below;

coupon payment=7.25% of 1000 =72.5 ; since semi annual coupons= 36.25 per half years;

we need to calculate IRR of cashflows and YTM=2*IRR (since payments are semi annual)

Bond (Annual payment)

Years 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18
Price 754.08

Coupon payment

36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25
Par value 1000
Total cashflows -754.08 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 36.25 1036.25
IRR 5.875%
YTM =2*IRR 11.750%

cost of debt=11.75%

---

WACC= (E/V)*ke +(D/V)*kd*(1-tax rate)

where E=market value Equity ;D=debt ;V=E+D =120,000,000 +80,000,000 =200000000

ke=cost of equity ;kd=cost of debt

WACC= (120,000,000/200000000)*17 +(80,000,000/200000000)*11.75*(1-.35) =13.255%